# PPA-Game: Characterizing and Learning Competitive Dynamics Among Online Content Creators

### v1

@@ Abstract
We introduce the Proportional Payoff Allocation Game (PPA-Game) to model how agents, akin to content creators on platforms like YouTube and TikTok, compete for divisible resources and consumers’ attention. Payoffs are allocated to agents based on heterogeneous weights, reflecting the diversity in content quality among creators. Our analysis reveals that although a pure Nash equilibrium (PNE) is not guaranteed in every scenario, it is commonly observed, with its absence being rare in our simulations. Beyond analyzing static payoffs, we further discuss the agents’ online learning about resource payoffs by integrating a multi-player multi-armed bandit framework. We propose an online algorithm facilitating each agent’s maximization of cumulative payoffs over $T$ rounds. Theoretically, we establish that the regret of any agent is bounded by $O(\log^{1+\eta}T)$ for any $\eta>0$. Empirical results further validate the effectiveness of our approach.
@@ 1 Introduction
Online recommender systems, exemplified by YouTube and TikTok, are now central to our daily digital experiences. From an economic viewpoint, these systems represent intricate ecosystems composed of multiple stakeholders: consumers, content creators, and the platform itself  (Boutilier et al., 2023; Ben-Porat et al., 2020). Although a significant portion of current research delves into consumer preferences and individualized recommendations, the role and behavior of content creators are also crucial  (Zhan et al., 2021; Boutilier et al., 2023). They not only shape the pool of recommended content but also influence the system’s stability, fairness, and long-term utility  (Ben-Porat and Tennenholtz, 2018; Patro et al., 2020; Boutilier et al., 2023; Mladenov et al., 2020). A major challenge when capturing the dynamics of content creators is characterizing the competitive exposure amongst them, given the finite demand for different content. Consequently, understanding the competitive dynamics among content creators becomes paramount.
@@
Game theory offers a useful framework to characterize competitive behaviors among multiple players  (Nisan et al., 2007). While numerous studies have employed game-theoretical analysis on content creators within online platforms, many  (Boutilier et al., 2023; Ben-Porat and Tennenholtz, 2018; Hron et al., 2022; Jagadeesan et al., 2022; Ben-Porat et al., 2019a) primarily explore Nash equilibrium properties, largely overlooking the dynamics of content creators and their convergence to the equilibrium. Some research attempts to bridge this gap, considering different allocation models for creators producing analogous content. Specifically, Ben-Porat et al. (2019b, 2020); Liu et al. (2020, 2021) posit that only top-tier content creators receive exposure, and Xu et al. (2023) proposes an averaged allocation rule. While these dynamics yield persuasive findings, they exhibit limitations stemming from the disconnect between their foundational principles and real-world contexts. For instance, top-winner dynamics may conflict with diversity principles in online recommendation platforms, while an averaged allocation rule could overlook the nuanced specializations that creators bring to diverse content domains.
@@
To bridge the gap between theoretical models and the complex dynamics observed among online content creators, we introduce the Proportional Payoff Allocation Game (PPA-Game) as a novel model to capture the competitive dynamics of online content creators. Adapting from  (Ben-Porat et al., 2019b, 2020; Prasad et al., 2023), we assume a game with $N$ players (content creators) and $K$ resources (content topics). In this context, players need to decide which resources to select. Each resource has an associated total payoff (the exposure for each content topic), and individual players have specific weights (quality metrics) assigned to each resource. When several players choose the same resource, its payoff is distributed proportionally according to player weights—a mechanism resonating with other game-theoretical frameworks  (Caragiannis and Voudouris, 2016; Hoefer and Jiamjitrak, 2017; Anbarci et al., 2023).
We first examine the existence of pure Nash equilibria (PNE) in our game model, uncovering that PNE may not always exist in certain scenarios. While this result could initially appear frustrating, our synthetic experiments suggest that instances lacking PNE are infrequent. Further investigation has identified several conditions under which the existence of PNE is typically ensured. These conditions, which have wide-ranging applicability in practical scenarios, ensure the presence of PNE if any of the following criteria are met: (1) The resources’ payoffs follow a long-tailed distribution, wherein the payoff ratio between any two resources either surpasses a specific constant $c_{0}$ or falls below its reciprocal $1/c_{0}$. (2) Players have the same weights across different resources. (3) Players have the same weights within each resource. (4) All resources have equal payoffs.
@@
Leveraging the PPA-Game, we introduce a Multi-player Multi-Armed Bandit (MPMAB) framework  (Rosenski et al., 2016; Besson and Kaufmann, 2018; Boursier and Perchet, 2019, 2020; Shi et al., 2020, 2021; Huang et al., 2022; Lugosi and Mehrabian, 2022) to simulate environments where content creators lack prior information about content demand and their quality in various topics. In this context, the ‘players’ represent content creators, while ‘arms’ denote content topics. This model entails players making simultaneous decisions over multiple rounds, with payoffs determined by the PPA-Game mechanism. Motivated by  (Xu et al., 2023; Boursier and Perchet, 2020), we assume that selfish players, motivated by their own interests, aim to maximize their own cumulative payoffs over $T$ rounds. Within this MPMAB structure, inspired by  (Bistritz and Leshem, 2018, 2021; Pradelski and Young, 2012), we delineate a novel online learning strategy for players and prove that each player’s regret is bounded by $O(\log^{1+\eta}T)$ for any $\eta>0$. Synthetic experiments validate the effectiveness of our approach.
@@ 2 Related Works
Game-theoretical Analysis for Content Creators.
While much of the research in this area has focused on consumer preferences and recommendations (Boutilier et al., 2023; Zhan et al., 2021), some studies have examined the competition among content creators using game-theoretical models (Boutilier et al., 2023; Ben-Porat and Tennenholtz, 2018; Hron et al., 2022; Jagadeesan et al., 2022; Ben-Porat et al., 2019a, b, 2020; Xu et al., 2023; Liu et al., 2020, 2021). A subset of these studies attempts to capture the dynamics among content creators. However, they often oversimplify exposure allocation, either favoring top-quality content creators (Ben-Porat et al., 2019b, 2020; Liu et al., 2020, 2021) or assuming equal exposure distribution (Xu et al., 2023). In contrast to these approaches, we assume that creators bring distinct expertise and excel in producing varied content types. Furthermore, different from Yao et al. (2023)’s study on coarse correlated equilibria with no-regret dynamics, our work distinctively focuses on pure Nash equilibria.
Proportional Allocation.
Proportional allocation is a well-explored mechanism in game theory, with various applications from resource distribution (Kelly, 1997; Caragiannis and Voudouris, 2016; Syrgkanis and Tardos, 2013; Johari, 2007) to Blotto games (Anbarci et al., 2023; Klumpp et al., 2019; Arad and Rubinstein, 2012; Friedman, 1958). These studies often focus on the strategic allocation of a single divisible resource or budget distribution across multiple contests. In contrast, our research diverges significantly by considering multiple divisible resources, with players needing to make strategic choices among these resources. This approach differs notably from the proportional coalition formation game explored by Hoefer and Jiamjitrak (2017), where payoffs within coalitions are divided based on player weights. Additionally, while the weighted congestion game (Bhawalkar et al., 2014) assumes that players have uniform weights across resources, our model accounts for players having different weights on various resources, adding a layer of complexity and realism to our game-theoretical analysis.
Multi-player Multi-Armed Bandits.
The Multi-Armed Bandit problem and its multi-player variant (MPMAB) have been extensively explored (Lai et al., 1985; Bubeck et al., 2012; Slivkins et al., 2019; Boursier and Perchet, 2022). In MPMAB, a common assumption is the absence of payoffs during player collisions (Boursier and Perchet, 2022). Various studies propose different payoff allocation mechanisms for collision scenarios, ranging from awarding the entire payoff to the highest bidder (Liu et al., 2020, 2021; Jagadeesan et al., 2021; Basu et al., 2021) to implementing thresholds or capacities for each arm (Bande and Veeravalli, 2019; Youssef et al., 2021; Bande et al., 2021; Magesh and Veeravalli, 2021; Wang et al., 2022a, b). Unlike the models that either maximize total welfare or offer reduced payoffs during collisions (Shi and Shen, 2021; Pacchiano et al., 2021; Boyarski et al., 2021), our study focuses on individual payoff maximization, highlighting player selfishness.
@@ 3.1 Background on Game Theory
A strategy profile $\bm{\pi}\in\mathcal{S}^{N}$ for a game defined by $\mathcal{G}=\left(\mathcal{S},\{\mathcal{U}_{j}\}_{j=1}^{N}\right)$ is termed an $\epsilon$-Nash equilibrium if, for every strategy $\pi^{\prime}\in\mathcal{S}$ and for each player $j\in[N]$, $\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})\leq\mathcal{U}_{j}(\bm{\pi})+\epsilon$. If $\bm{\pi}$ achieves a 00-Nash equilibrium, it’s termed a pure Nash equilibrium (PNE).
@@ 3.2 Game Formulation
In modern recommendation platforms like YouTube and TikTok, content creators compete for viewer attention by producing content on diverse topics. In this context, content creators serve as the ‘players,’ while the various topics they generate act as the ‘resources.’ The ’payoff’ associated with each resource represents the overall attention or exposure a topic receives. Since creators possess varying expertise across different topics, the exposure tends to be proportionally distributed based on their respective influence or proficiency in that topic.
Several models closely relate to our formulation, particularly the studies by (Bhawalkar et al., 2014; Ben-Porat et al., 2019b, 2020; Liu et al., 2020, 2021; Xu et al., 2023). However, models from (Ben-Porat et al., 2019b, 2020; Liu et al., 2020, 2021) assume that only the player $j$ with the highest weight $w_{j,k}$ for a resource $k$ receives the payoff. Xu et al. (2023) assumed an average payoff allocation to players choosing the resource. The weighted congestion game (Bhawalkar et al., 2014) assumes that players have the same weights on all resources. Hence, we note that all these formulations diverge from realistic scenarios of online recommender systems, as explained above.
@@ 3.3 Analysis on Pure Nash Equilibrium (PNE)
In this subsection, we investigate Pure Nash Equilibria (PNEs) due to their role as the most stringent equilibria in game theory (Nisan et al., 2007) and their importance for the stability of platforms (Ben-Porat and Tennenholtz, 2018). This exploration is consistent with previous research (Ben-Porat et al., 2020; Xu et al., 2023; Hron et al., 2022) that has examined PNEs within recommender systems.
@@ 3.3.1 High Probability of PNE Existence
We first acknowledge that PNE may not always exist in certain scenarios (see Appendix A.1 for detailed illustrative example):
@@
While this result might seem frustrating, subsequent synthetic experiments largely demonstrate the typical existence of PNEs in most cases in a large-scale simulation study (see Appendix A.2 for detailed experimental setup):
@@
We randomly choose $N$ and $K$ from the range $[10,100]$ and sample resources’ payoffs and players’ weights from four distributions supported on $[0,1]$. Among the $20,000,000$ randomly generated game configurations we tested, PNE was absent in only one instance.

### v2

@@ Abstract
In this paper, we present the Proportional Payoff Allocation Game (PPA-Game), which characterizes situations where agents compete for divisible resources. In the PPA-game, agents select from available resources, and their payoffs are proportionately determined based on heterogeneous weights attributed to them. Such dynamics simulate content creators on online recommender systems like YouTube and TikTok, who compete for finite consumer attention, with content exposure reliant on inherent and distinct quality. We first conduct a game-theoretical analysis of the PPA-Game. While the PPA-Game does not always guarantee the existence of a pure Nash equilibrium (PNE), we identify prevalent scenarios ensuring its existence. Simulated experiments further prove that the cases where PNE does not exist rarely happen. Beyond analyzing static payoffs, we further discuss the agents’ online learning about resource payoffs by integrating a multi-player multi-armed bandit framework. We propose an online algorithm facilitating each agent’s maximization of cumulative payoffs over $T$ rounds. Theoretically, we establish that the regret of any agent is bounded by $O(\log^{1+\eta}T)$ for any $\eta>0$. Empirical results further validate the effectiveness of our online learning approach.
@@ 1 Introduction
Online recommender systems, exemplified by platforms such as YouTube and TikTok, are now central to our daily digital experiences. From an economic viewpoint, these systems represent intricate ecosystems composed of multiple stakeholders: consumers, content creators, and the platform itself  (Boutilier et al., 2023; Ben-Porat et al., 2020). Although a significant portion of current research delves into consumer preferences and individualized recommendations, the role and behavior of content creators are also crucial  (Zhan et al., 2021; Boutilier et al., 2023). They not only shape the pool of recommended content but also influence the system’s stability, fairness, and long-term utility  (Ben-Porat and Tennenholtz, 2018; Patro et al., 2020; Boutilier et al., 2023; Mladenov et al., 2020). A major challenge when capturing the dynamics of content creators is characterizing the competitive exposure among them, given the finite demand for different content. Consequently, understanding the competitive dynamics among content creators becomes paramount.
@@
Game theory offers a useful framework to characterize competitive behaviors among multiple players  (Nisan et al., 2007). While numerous studies have employed game-theoretical analysis on content creators within online platforms, many  (Yao et al., 2024b, a, 2023; Ben-Porat and Tennenholtz, 2018; Hron et al., 2022; Jagadeesan et al., 2022; Ben-Porat et al., 2019a) primarily explore Nash equilibrium properties, largely overlooking the dynamics of content creators and their convergence to the equilibrium. Some research attempts to bridge this gap, considering different allocation models for creators producing analogous content  (Ben-Porat et al., 2019b, 2020; Xu et al., 2023; Liu et al., 2020, 2021). For instance, Ben-Porat et al. (2019b, 2020); Liu et al. (2020, 2021) assumed that only top-tier content creators receive exposure, and Xu et al. (2023) proposed an averaged allocation rule. While these dynamics offer valuable insights, they also have limitations due to the gap between their modeling assumptions and real-world contexts. For instance, top-winner dynamics (Ben-Porat et al., 2019b, 2020; Liu et al., 2020, 2021) may conflict with diversity principles in online recommendation platforms, while an averaged allocation rule (Xu et al., 2023) could overlook the nuanced specializations that creators bring to diverse content domains.
@@
To bridge the gap between theoretical models and the complex dynamics observed among online content creators, we introduce the Proportional Payoff Allocation Game (PPA-Game) as a novel model to capture the competitive dynamics of online content creators. Adapting from  (Ben-Porat et al., 2019b, 2020; Prasad et al., 2023), we assume a game with $N$ players (content creators) and $K$ resources (content topics). In this context, players need to decide which resources to select. Each resource has an associated total payoff (indicative of the exposure for each content topic), and individual players have specific weights (or quality metrics) assigned to each resource. When several players choose the same resource, its payoff is distributed proportionally according to player weights—a mechanism resonating with other game-theoretical frameworks  (Caragiannis and Voudouris, 2016; Kelly, 1997; Hoefer and Jiamjitrak, 2017; Anbarci et al., 2023).
We first explore the existence of the pure Nash equilibria (PNE) within the game, revealing that PNE might not exist in particular scenarios. We further identify several prevalent conditions where PNE’s existence is guaranteed. These conditions, with broad applicability in real-world scenarios, guarantee PNE existence if any of the following holds: (1) The payoffs of the resources exhibit a long-tailed distribution, meaning the ratio between the payoffs of any two resources either exceeds a specific constant $c_{0}>1$ or is less than its reciprocal $1/c_{0}$. (2) All players maintain the same weights across different resources. (3) Every resource has identical payoffs. To further investigate the probability of PNE’s existence, we conducted simulation experiments. Our results indicate that such situations where PNE is absent rarely occur.
@@
Leveraging the PPA-Game, we introduce a Multi-player Multi-Armed Bandit (MPMAB) framework  (Rosenski et al., 2016; Besson and Kaufmann, 2018; Boursier and Perchet, 2019, 2020; Shi et al., 2020, 2021; Huang et al., 2022; Lugosi and Mehrabian, 2022) to simulate environments where content creators lack prior information about content demand and their quality in various topics. In this context, the ‘players’ represent content creators, while ‘arms’ denote content topics. This model entails players making simultaneous decisions over multiple rounds, with payoffs determined by the PPA-Game mechanism. Inspired by Xu et al. (2023); Boursier and Perchet (2020), we assume that selfish players, motivated by their own interests, aim to maximize their own cumulative payoffs over $T$ rounds. Within this MPMAB structure, inspired by  (Bistritz and Leshem, 2018, 2021; Pradelski and Young, 2012), we delineate a novel online learning strategy for players and prove that each player’s regret is bounded by $O(\log^{1+\eta}T)$ for any $\eta>0$. Synthetic experiments validate the effectiveness of our approach.
In summary, our key contributions include:
  • We introduce the novel Proportional Payoff Allocation Game (PPA-Game) to characterize competitive dynamics among online content creators.
  • We comprehensively examine the PNE’s existence within the PPA-Game, identifying broad scenarios where the existence of PNE is prevalent, as proved by simulation results.
  • Building on the PPA-Game, we introduce a new Multi-player Multi-Armed Bandit (MPMAB) paradigm for simulating real-time decision-making by content creators.
  • We propose a novel online algorithm, with bounded regret, for modeling player competition. Empirical results demonstrate its efficacy.
The remainder of this paper is structured as follows. Section 2 provides a systematic review of related work. Section 3 introduces the PPA-Game, examines its theoretical properties, and characterizes the conditions for PNE existence. In Section 4, we present the MPMAB framework, propose an online learning algorithm, and establish its theoretical guarantees. Section 5 reports synthetic experiment results, and Section 6 summarizes our contributions.
@@ 2 Related Works
Game-theoretical Analysis for Content Creators
While a significant portion of existing research concentrates on consumer preferences and individualized recommendations (Boutilier et al., 2023; Zhan et al., 2021), a few delve into the competitive dynamics among content creators, framing this competition through game-theoretical models (Yao et al., 2024b, a; Boutilier et al., 2023; Ben-Porat and Tennenholtz, 2018; Hron et al., 2022; Jagadeesan et al., 2022; Ben-Porat et al., 2019a, b, 2020; Xu et al., 2023; Liu et al., 2020, 2021; Yao et al., 2023). For instance, works by Hron et al. (2022); Jagadeesan et al. (2022) model content creator strategies as vectors within a designated space, offering insights into various Nash equilibrium properties. Meanwhile, Ben-Porat et al. (2019a) conceptualized this competition using facility location games (Anderson et al., 1992), though they did not incorporate the learning dynamics inherent to content creators.
A subset of these studies attempts to capture the dynamics among content creators (Ben-Porat et al., 2019b, 2020; Liu et al., 2020, 2021; Xu et al., 2023). Specifically, studies like Ben-Porat et al. (2019b, 2020); Liu et al. (2020, 2021) assume that only top-quality content creators receive exposure for each content topic. This perspective seems overly simplistic, especially considering the platforms’ need to equitably distribute exposure to retain a diverse array of creators. On the other hand, Xu et al. (2023) adopted an averaging allocation rule, suggesting that creators share content topics’ exposure equally on average. However, this fails to account for the real-world scenarios where creators bring distinct expertise and excel in producing varied content types.
Proportional Allocation
Proportional allocation serves as a frequently adopted mechanism in game theory (Kelly, 1997; Caragiannis and Voudouris, 2016; Syrgkanis and Tardos, 2013; Johari, 2007; Anbarci et al., 2023; Hoefer and Jiamjitrak, 2017). However, the game formulations in these works vary considerably from our formulation. Works focusing on the proportional allocation mechanism (Kelly, 1997; Caragiannis and Voudouris, 2016; Syrgkanis and Tardos, 2013; Johari, 2007) delves into the distribution of a singular, infinitely divisible resource amongst multiple competing players. In these models, players strategically choose the resource bid values. In contrast, studies like Anbarci et al. (2023); Klumpp et al. (2019) navigate the Blotto games (Arad and Rubinstein, 2012; Friedman, 1958), where players allocate budgets across various battlefields or contests. Here, the objective hinges on amplifying total winnings, with victory rates in battles proportional to a player’s allocated budget. Distinct from these models, our work focuses on multiple divisible resources and players need to choose among different resources.
Furthermore, Hoefer and Jiamjitrak (2017) explored a proportional coalition formation game, where players create coalitions (i.e., player groups). Within this framework, coalition payoffs get proportionally divided based on player weights. By contrast, our approach considers the distribution of several divisible resources where identical player groups may gain varying payoffs from different resources.
Multi-player Multi-Armed Bandits
The Multi-Armed Bandit (MAB) problem (Lai et al., 1985; Bubeck et al., 2012; Slivkins et al., 2019; Lattimore and Szepesvári, 2020) has widely studied across various settings (Kveton et al., 2015; Lattimore et al., 2016; Agrawal and Devanur, 2016; Xu et al., 2022; Ferreira et al., 2022). An extension to the classic MAB is the multi-player multi-armed bandit (MPMAB) (Liu and Zhao, 2008; Jouini et al., 2009, 2010; Brânzei and Peres, 2021), which studies scenarios wherein multiple players interact within a single MAB instance (Boursier and Perchet, 2022). Notably, a majority of MPMAB research posits that players receive no payoff in the event of a collision (Rosenski et al., 2016; Besson and Kaufmann, 2018; Boursier and Perchet, 2019, 2020; Shi et al., 2020, 2021; Huang et al., 2022; Lugosi and Mehrabian, 2022), as comprehensively surveyed in (Boursier and Perchet, 2022).
Several studies have proposed diverse payoff allocation mechanisms within the MPMAB framework when collisions occur. For instance, works by Liu et al. (2020, 2021); Jagadeesan et al. (2021); Basu et al. (2021) assume that the player with the highest expected payoff get the entire payoff in such scenarios. Others have introduced concepts like thresholds (Bande and Veeravalli, 2019; Youssef et al., 2021; Bande et al., 2021; Magesh and Veeravalli, 2021) or capacities (Wang et al., 2022a, b) for each arm. In addition, Shi and Shen (2021); Pacchiano et al. (2021) assumed reduced payoffs for each player during collisions. Boyarski et al. (2021) presented a model where payoffs are heterogeneous, yet they do not ensure that the sum of payoffs for collided players aligns with the arm’s total payoff. Crucially, while many of these studies emphasize solutions that maximize total welfare, we make the assumption of player selfishness, focusing on individual payoff maximization. Lastly, although Xu et al. (2023) also operated within the MPMAB framework, they advocate an averaging allocation mechanism, a departure from real-world applicability and distinct from our formulation.
@@ 3.1 Background on Game Theory
A strategy profile $\bm{\pi}\in\mathcal{S}^{N}$ for a game defined by $\mathcal{G}=\left(\mathcal{S},\{\mathcal{U}_{j}\}_{j=1}^{N}\right)$ is termed an $\epsilon$-Nash equilibrium if, for every strategy $\pi^{\prime}\in\mathcal{S}$ and for each player $j\in[N]$, $\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})\leq\mathcal{U}_{j}(\bm{\pi})+\epsilon$. If $\bm{\pi}$ achieves a 00-Nash equilibrium, it’s termed a pure Nash equilibrium.
We focus on PNE because they represent the most robust form of equilibrium in game theory (Nisan et al., 2007) and are essential for ensuring platform stability (Ben-Porat and Tennenholtz, 2018). Previous studies in recommender systems have also emphasized the importance of PNE in understanding system behavior and fairness (Ben-Porat et al., 2020; Xu et al., 2023; Hron et al., 2022). Furthermore, in the context of our proposed game in Section 3.2, we note that since it is finite, Mixed Nash Equilibria (MNE) always exist (Nisan et al., 2007). However, focusing on PNE is crucial for capturing the most stringent, predictable, and enforceable outcomes in practice.
@@ 3.2 Game Formulation
In modern recommendation platforms like YouTube and TikTok, content creators compete for viewer attention by producing content on a wide range of topics. In this setting, the content creators are the ‘players,’ while the topics they generate act as the ‘resources.’ The ‘payoff’ associated with each resource reflects the overall attention or exposure a topic receives. Since creators possess varying levels of expertise across different topics, exposure is typically distributed in proportion to their influence or proficiency in each specific topic.
Several models are closely related to our formulation, particularly those by Bhawalkar et al. (2014); Ben-Porat et al. (2019b, 2020); Liu et al. (2020, 2021); Xu et al. (2023). However, the models in (Ben-Porat et al., 2019b, 2020; Liu et al., 2020, 2021) assume that only the player $j$ with the highest weight $w_{j,k}$ for a resource $k$ receives the payoff. In contrast, Xu et al. (2023) assume an average payoff allocation for players choosing the same resource. The weighted congestion game model (Bhawalkar et al., 2014) assumes that players have identical weights across all resources. Therefore, we argue that these formulations diverge from the more realistic dynamics of online recommender systems, as outlined above.
@@ 3.3 Analysis on Pure Nash Equilibrium (PNE)
In this subsection, we investigate properties of Pure Nash Equilibria.
@@ 3.3.1 High Probability of PNE Existence
We first acknowledge that PNE may not always exist in certain scenarios (see Section A.1 for a detailed illustrative example):
@@
While this result may seem discouraging, large-scale simulations suggest that PNEs typically exist in most cases. Our synthetic experiments support this finding, as detailed in Section A.2.
@@
We randomly select $N$ and $K$ from the range $[10,100]$ and sample resource payoffs and player weights from four distributions supported on $[0,1]$. Among $20,000,000$ randomly generated game configurations, a PNE fails to exist in only one instance.